OpenStudy (maheshmeghwal9):

Show that f(x) is continuous at x=0:--

OpenStudy (maheshmeghwal9):

f(x)={xsin*1/x; x\[\neq\]0} & 0 for x=0.

OpenStudy (experimentx):

\[ \lim_{x \rightarrow 0} x \sin(1/x) = 0 * some \; value \; between \; -1\; and\; +1 = 0\]

OpenStudy (maheshmeghwal9):

k!

OpenStudy (anonymous):

Let \(\delta=\varepsilon\). Then\[|x-0|=|x|<\delta,\]\[|f(x)-f(0)|=|f(x)|=\left|x\sin\left(\frac{1}{x}\right)\right|=|x|\left|\sin\left(\frac{1}{x}\right)\right|\leqslant|x|<\delta=\varepsilon.\]Hence, \(f\) is continuous at \(0\).

OpenStudy (maheshmeghwal9):

please tell me about writing left hand limit & right hand limit for f(x) given

OpenStudy (experimentx):

x->a+ means it approaches from right side, in our case, x>0 x->a- means it approaches from right side, if your case, x<0 Either cases, our limit would be zero ..

OpenStudy (maheshmeghwal9):

Thanx to all!